OpenStudy (anonymous):

Using a lever, a person applies 60 N of force, and this action lifts a 30 kg rock at the other end. What is the mechanical advantage of his action?

OpenStudy (anonymous):

he would need F=mg=30*9.8=around 300N to keep lift it with his hands (against gravity)

OpenStudy (anonymous):

Mechanical advantage is the difference?

OpenStudy (anonymous):

well the difference in mechanical energy that could be spent lifting with hands vs lever cant be found with the data at hand so i would say that the mechanical advantage is that with a force 240 N less he can lift it with lever

OpenStudy (anonymous):

well 294-60=234 to be exact